# Kernel PCA for Out-of-Distribution Detection

### v1

@@ I Details of OoD Detectors Involved in Comparison Experiments
We follow the settings in KNN [14] and include CSI [23] and SSD [24] into the comparisons in Sec.4.1. The 2 methods adopt the contrastive losses to train DNNs. In the comparison results of Tab.1 and the following Tab.II.1, the reported detection results w.r.t. CSI and SSD are directly from [7], and are obtained by executing the Mahalanobis detector on learned features of DNNs trained by CSI and SSD. Refer to [14] for more details.
II Supplementary Experiment Results
Table II.1: The detection performance of different methods (ResNet18 trained on CIFAR10) on each OoD data set.
method OoD data sets
SVHN LSUN iSUN Textures Places365 AVERAGE
@@ II Supplementary Experiment Results
Without Supervised Contrastive Learning
MSP 59.66 91.25 45.21 93.80 54.57 92.12 66.45 88.50 62.46 88.64 57.67 90.86
ODIN 20.93 95.55 7.26 98.53 33.17 94.65 56.40 86.21 63.04 86.57 36.16 92.30
Energy 54.41 91.22 10.19 98.05 27.52 95.59 55.23 89.37 42.77 91.02 38.02 93.05
GODIN 15.51 96.60 4.90 99.07 34.03 94.94 46.91 89.69 62.63 87.31 32.80 93.52
Mahalanobis 9.24 97.80 67.73 73.61 6.02 98.63 23.21 92.91 83.50 69.56 37.94 86.50
KNN 24.53 95.96 25.29 95.69 25.55 95.26 27.57 94.71 50.90 89.14 30.77 94.15
@@
With Supervised Contrastive Learning
CSI 37.38 94.69 5.88 98.86 10.36 98.01 28.85 94.87 38.31 93.04 24.16 95.89
SSD 1.51 99.68 6.09 98.48 33.60 95.16 12.98 97.70 28.41 94.72 16.52 97.15
KNN 2.42 99.52 1.78 99.48 20.06 96.74 8.09 98.56 23.02 95.36 11.07 97.93
@@
II.1 Comparisons with Nearest Neighbor Searching
Tab.II.1 illustrates the comparison results on the CIFAR10 benchmark between our CoP/CoRP and the KNN detector [7]. Similar as the comparisons on the ImageNet-1K benchmark in Tab.1, CoRP outperforms other baselines in both the standard training and the supervised contrastive learning with lower FPR and higher AUROC average values.
Table II.2: Comparisons on the inference time of the KNN detector [7] and our CoP and CoRP. The ImageNet data is of the in-distribution. The experiments are executed on the same machine for a fair comparison. The nearest neighbor searching of KNN is implemented via Faiss [45].
method time complexity inference time (ms, per sample)
KNN $\mathcal{O}(N_{\rm tr})$approx 15.59
CoP and CoRP (ours) $\mathcal{O}(1)$approx 0.06
Table II.3: Comparisons on the detection performance between the regularized reconstruction error [8] and the proposed CoP and CoRP fused with other OoD scores (MSP, Energy, ReAct and BATS) on each OoD data set (MobileNet pre-trained on ImageNet-1K).
method OoD data sets
iNaturalist SUN Places Textures AVERAGE
@@ II.1 Comparisons with Nearest Neighbor Searching
ODIN 58.54 87.51 57.00 85.83 59.87 84.77 52.07 85.04 56.87 85.79
Mahalanobis 62.11 81.00 47.82 83.66 52.09 83.63 92.38 33.06 63.60 71.01
ViM 91.83 77.47 94.34 70.24 93.97 68.26 37.62 92.65 79.44 77.15
DICE 43.28 90.79 38.86 90.41 53.48 85.67 33.14 91.26 42.19 89.53
@@
Tab.II.2 shows the time-consuming comparison between ours and the KNN detector during inference. The ImageNet-1K benchmark consists of 14,197,122 training images. KNN needs to store and iterate all these features to search the nearest neighbor and costs around 15.59 ms for a new sample, while the iterating process is omitted in our method, resulting in a much higher processing speed, less than 1 ms per sample.
II.2 Comparisons with Regularized Reconstruction Errors
Tab.II.3 shows the comparisons on the MobileNet between our KPCA-based reconstruction errors and the regularized reconstruction errors [8]. Both methods are evaluated under the same fusion strategy with other 4 OoD scores (MSP, Energy, ReAct and BATS). Similar as the results on ResNet50 in Tab.2, the proposed KPCA-based CoP and CoRP outperform [8] in almost all the cases, and CoRP achieves new SOTA OoD detection performance when fused with the ReAct detector. Such results again indicate that the introduced non-linearity of our KPCA could further benefit separating out-distribution samples from in-distribution ones.
@@ II.2 Comparisons with Regularized Reconstruction Errors
Figure 5: A sensitivity analysis on the explained variance ratio of CoRP. The average detection FPR and AUROC values of OoD data sets in CIFAR10 and ImageNet-1K benchmarks are reported. The Gaussian kernel width $\gamma$ and the dimension $M$ of RFFs are fixed.
@@
II.3 Sensitivity Analysis on CoRP
In this section, a detailed sensitivity analysis on the hyper-parameters of CoRP is provided. Recall the hyper-parameters in CoRP: the Gaussian kernel width $\gamma$, the dimension $M$ of RFFs and the number of columns $q$ of the dimensionality-reduction matrix $\boldsymbol{U}^{\Phi}_{q}$. In the following, we discuss the influence of each hyper-parameter, and report experiment results of the detection performance by varying one hyper-parameter with the other two fixed.
Effect of $q$
As discussed on CoP in Sec.4.3, $q$ determines the fraction of principal components preserved by PCA and is selected based on the explained variance ratio: A high explained variance ratio implies a large value of $q$. Fig.5 shows the detection FPR95 and AUROC values of CoRP w.r.t. different explained variance ratios on CIFAR10 and ImageNet-1K benchmarks. Similar as the case of CoP, an appropriate value of explained variance ratio needs to be carefully selected to achieve the optimal detection performance. Besides, a sufficiently large value of the explained variance ratio is sometimes no longer essential, which might be due to that the 2 concatenated kernels make the useful information for distinguishing out-distribution samples more concentrated in less principal components.
@@ II.3 Sensitivity Analysis on CoRP
The Gaussian kernel width $\gamma$ directly affects the mapped data distribution. For a large $\gamma$, then $k_{\rm gau}(\boldsymbol{z}_{1},\boldsymbol{z}_{2})=e^{-\gamma\left\|\boldsymbol{z}_{1}-\boldsymbol{z}_{2}\right\|_{2}^{2}}\approx 0$ for $\boldsymbol{z}_{1}\neq\boldsymbol{z}_{2}$, which indicates that the mapped features of $\boldsymbol{z}_{1}$ and $\boldsymbol{z}_{2}$ are (nearly) mutually-orthogonal. In this case, a PCA would become meaningless. For a small $\gamma$, then the KPCA-based reconstruction errors will approach the standard PCA-based ones, shown by [46]. Fig.6 illustrates the detection FPR95 and AUROC values of CoRP w.r.t. varied Gaussian kernel width $\gamma$ on CIFAR10 and ImageNet-1K benchmarks. Clearly, neither a too large nor a too small kernel width benefits the detection performance, and a mild value of $\gamma$ should be carefully tuned for different in-distribution data.
@@
The dimension $M$ of RFFs determines the approximation ability of RFFs towards the Gaussian kernel. As proved in [27], the larger the $M$, the better the RFFs approximate $k_{\rm gau}$. Fig.7 indicates the detection FPR95 and AUROC values of CoRP w.r.t. multiple values of the RFFs dimension $M$ on CIFAR10 and ImageNet-1K benchmarks. As $M$ increases, the detection performance gets improved since the RFFs better converge to the Gaussian kernel. Considering the computation efficiency of eigendecomposition on the covariance matrix of $\mathbb{R}^{M\times M}$, in the comparison experiments, we adopt $M=4m$ on CIFAR10 with $m=512$ for ResNet18, and $M=2m$ on ImageNet-1K with $m=2048$ for ResNet50 and $m=1280$ for MobileNet.
Figure 7: Sensitivity analysis on the dimension $M$ of RFFs of CoRP. The average detection FPR and AUROC values of OoD data sets in CIFAR10 and ImageNet-1K benchmarks are reported. The explained variance ratio and the Gaussian kernel width $\gamma$ are fixed.
III Supplementary Theoretical Results
@@
Recall $\boldsymbol{z}\in\mathbb{R}^{m}$ and suppose $\Phi:\mathbb{R}^{m}\rightarrow\mathbb{R}^{M}$ and $\boldsymbol{U}^{\Phi}\in\mathbb{R}^{M\times M}$ is the eigenvector matrix of the covariance matrix of the training data with $\boldsymbol{U}^{\Phi}=\left[\boldsymbol{U}^{\Phi}_{q},\boldsymbol{U}^{\Phi}_{p}\right]$ and $q+p=M$. For the reconstruction error $e^{\Phi}(\boldsymbol{\hat{z}})$ w.r.t. a new test sample $\boldsymbol{\hat{z}}\in\mathbb{R}^{m}$ in the mapped $\Phi(\boldsymbol{z})$-space, we have:
$\left(e^{\Phi}(\boldsymbol{\hat{z}})\right)^{2}$ $=\left\|(\Phi(\boldsymbol{\hat{z}})-\boldsymbol{\mu}^{\Phi})-\boldsymbol{U}^{\Phi}_{q}\boldsymbol{U}_{q}^{\Phi\top}(\Phi(\boldsymbol{\hat{z}})-\boldsymbol{\mu}^{\Phi})\right\|_{2}^{2}$ (III.1)

### v3

@@ Appendix A Details of representative OoD detectors
NNGuide [24] exploits the guidance of the Energy score on logits to boost the detection performance of the nearest neighbor search on features. Specifically, the training features are firstly scaled by their corresponding Energy scores, then KNN is executed on such re-scaled features for the new sample. The final detection score is set as the multiplication of the searched distance and the Energy score. The time and memory complexity of NNGuide is still the same $\mathcal{O}(N_{\rm tr})$ as that of KNN [7].
DML [28] decouples the maximum logits into two parts: the maximum cosine similarity (MaxCosine) and the maximum norm (MaxNorm), and employs their ensemble as the detection score. DML reveals that the cosine similarity and the feature norm jointly contribute to the effectiveness of the previous MSP [15] and MaxLogit [14] methods and designs new training losses from the perspective of feature collapse [55], so as to further improve the performance of MaxCosine and MaxNorm.
ASH [20] is a post-hoc detection method that removes the abnormal information in features. ASH includes two stages: removing a large portion of the features, and adjusting the remaining feature values by scaling up or assigning a constant value. ASH exhibits advantages over the classic ReAct method [18]: no global thresholds and stronger flexibility, and shows better detection performance.
SCALE [21] analyzes the rectification and scaling components of the ASH method, and improves ASH by only a scaling process in the post-hoc stage.
DDCS [22] investigates the effects of different channels based on existing feature-clipping detection methods, and proposes a channel-level anomalous activations pre-rectifying module so as to clip features more carefully for better detection performance.
We follow the settings in KNN [7] and include CSI [32] and SSD [33] into the comparisons in Section 5.1. The 2 methods adopt the contrastive losses to train DNNs. In the comparison results of Table 1 and the following Table 4, the reported detection results of CSI and SSD are directly from [7], and are obtained by executing the Mahalanobis detector on learned features of DNNs trained by CSI and SSD. Refer to [7] for more details.
Appendix B Supplementary experiment results on OoD detection
Table 4: The detection performance of different methods (ResNet18 trained on CIFAR10).
method OoD data sets AVERAGE
SVHN LSUN iSUN Textures Places365
@@ Appendix B Supplementary experiment results on OoD detection
Standard Training
MSP [15] 59.66 91.25 45.21 93.80 54.57 92.12 66.45 88.50 62.46 88.64 57.67 90.86
ODIN [16] 20.93 95.55 7.26 98.53 33.17 94.65 56.40 86.21 63.04 86.57 36.16 92.30
Energy [5] 54.41 91.22 10.19 98.05 27.52 95.59 55.23 89.37 42.77 91.02 38.02 93.05
GODIN [27] 15.51 96.60 4.90 99.07 34.03 94.94 46.91 89.69 62.63 87.31 32.80 93.52
Mahala [23] 9.24 97.80 67.73 73.61 6.02 98.63 23.21 92.91 83.50 69.56 37.94 86.50
KNN [7] 24.53 95.96 25.29 95.69 25.55 95.26 27.57 94.71 50.90 89.14 30.77 94.15
@@
Supervised Contrastive Learning
CSI [32] 37.38 94.69 5.88 98.86 10.36 98.01 28.85 94.87 38.31 93.04 24.16 95.89
SSD [33] 1.51 99.68 6.09 98.48 33.60 95.16 12.98 97.70 28.41 94.72 16.52 97.15
KNN [7] 2.42 99.52 1.78 99.48 20.06 96.74 8.09 98.56 23.02 95.36 11.07 97.93
@@
Table 5: Comparisons on the detection performance between the regularized reconstruction error [8] and our CoP and CoRP fused with other OoD scores (MSP, Energy, ReAct and BATS) on each OoD data set (MobileNet trained on ImageNet-1K). Best average results are highlighted with underlines.
method OoD data sets AVERAGE
iNaturalist SUN Places Textures
@@
ODIN [16] 58.54 87.51 57.00 85.83 59.87 84.77 52.07 85.04 56.87 85.79
Mahala [23] 62.11 81.00 47.82 83.66 52.09 83.63 92.38 33.06 63.60 71.01
ViM [35] 91.83 77.47 94.34 70.24 93.97 68.26 37.62 92.65 79.44 77.15
DICE [34] 43.28 90.79 38.86 90.41 53.48 85.67 33.14 91.26 42.19 89.53
@@
NNGuide [24] 45.73 91.19 51.03 87.87 60.60 84.44 29.50 92.47 46.72 88.99
ASH-B [20] 31.46 94.28 38.45 91.61 51.80 87.56 20.92 95.07 35.66 92.13
ASH-S [20] 39.10 91.94 43.62 90.02 58.84 84.73 13.12 97.10 38.67 90.95
Table 4 illustrates the comparison results on the CIFAR10 benchmark between our CoP/CoRP and the KNN detector [7]. Similar as the comparisons on the ImageNet-1K benchmark in Table 1, CoRP outperforms other baselines in both the standard training and the supervised contrastive learning with lower FPR and higher AUROC average values.
Table 5 shows the comparison results on the ImageNet-1K benchmark between our CoP/CoRP and the regularized reconstruction error [8] of MobileNet [51]. Similar as the case of ResNet50 in Table 3, under the same fusion trick with other detection scores, our KPCA reconstruction errors of CoP and CoRP significantly outperform the regularized PCA reconstruction error of [8], implying the substantial improvements by characterizing the non-linear data distribution of the InD and OoD features via the devised two proper non-linear kernels.
Appendix C Analytical discussions on kernels
In this section, more in-depth discussions are drawn on the effects of cosine and Gaussian kernels in CoP and CoRP for OoD detection in Section C.1 and Section C.2, respectively. A comprehensive sensitivity analysis on the involved hyper-parameters in CoP and CoRP are presented in Section C.3.
C.1 Effects of the cosine kernel
The cosine kernel in CoP and CoRP appears an indispensable basis in alleviating the linear inseparability in InD and OoD features. The reason for its effectiveness lies in the imbalanced feature norms $\|\boldsymbol{z}\|_{2}$ between InD and OoD features, which has also been observed in preceding works [7, 6, 32]. Figure 3 shows the feature norms of multiple InD and OoD data sets, from which one can find clear disparities of the InD and OoD feature norms. The cosine kernel $k_{\rm cos}$ and the corresponding feature mapping $\phi_{\rm cos}$ in Equation 6 thereby normalize the feature norms and facilitate the separability between InD and OoD data.
Figure 4 illustrates the t-SNE visualization on the InD and OoD features to further imply the importance of the cosine kernel. As shown in the middle panel of Figure 4, the Gaussian kernel alone fails in creating separable InD and OoD features in the mapped space and actually leads to a complete mess of the mapped features. In contrast, the cosine kernel significantly alleviates the linearly-inseparability of features, shown in the right panel of Figure 4.
Ablation studies on the cosine feature mapping $\phi_{\rm cos}$ in CoP and CoRP are executed to verify its indispensability for OoD detection. Specifically, CoP without $\phi_{\rm cos}$ reduces to a standard PCA on the $\boldsymbol{z}$-space, and CoRP without $\phi_{\rm cos}$ reduces to KPCA with a Gaussian kernel. Table 6 shows the corresponding detection FPR and AUROC values on each OoD data set in the ImageNet-1K benchmark of ablations on $\phi_{\rm cos}$. Both the standard PCA and the Gaussian KPCA exhibit worse detection performance than CoP and CoRP. Particularly, the single Gaussian kernel in KPCA actually results in a complete failure in detecting OoD samples with nearly 95% FPR values. Therefore, the cosine kernel is essential in characterizing the non-linearity in InD and OoD features and critical for the superior performance of the KPCA detector.
Figure 3: A density histogram of the imbalanced norms of InD and OoD features. InD: CIFAR10 and ImageNet-1K. OoD: LSUN and places365, SUN and Textures.
Figure 4: T-SNE visualization of the original features (left), mapped features w.r.t a Gaussian kernel (middle) and mapped features w.r.t a cosine kernel (right).
C.2 Alternative kernels
The design of the Gaussian kernel in CoRP is motivated by the useful $\ell_{2}$ distance on $\frac{\boldsymbol{z}}{\|\boldsymbol{z}\|_{2}}$ in the KNN detector [7]. The Gaussian kernel preserves the beneficial $\ell_{2}$ distance relationship in $\phi_{\rm cos}(\boldsymbol{z})$-space through the RFFs mapping. In this section, we provide two alternative choices aside from the cosine-Gaussian kernel.
  • The cosine-Laplacian kernel explores the $\ell_{1}$ distance in $\phi_{\rm cos}(\boldsymbol{z})$-space via the Laplacian kernel $k_{\rm lap}$:
    $k_{\rm lap}(\boldsymbol{z}_{1},\boldsymbol{z}_{2})=e^{-\gamma\|\boldsymbol{z}_{1}-\boldsymbol{z}_{2}\|_{1}}.$ (16)
    To construct the RFFs for $k_{\rm lap}$, the sampling distribution of the $\omega_{i}$ in Equation 4 is a Cauchy distribution $\omega\sim p(\omega)=\frac{\gamma^{2}}{\pi\gamma(\omega^{2}+\gamma^{2})}$.
  • The cosine-polynomial kernel does not hold the $\ell_{1}$ nor $\ell_{2}$ distance-preserving property for $\phi_{\rm cos}(\boldsymbol{z})$-space, as the polynomial kernel is defined as:
    $k_{\rm poly}(\boldsymbol{z}_{1},\boldsymbol{z}_{2})=(\boldsymbol{z}_{1}^{\top}\boldsymbol{z}_{2}+c)^{d}.$ (17)
    To obtain an explicit feature mapping for $k_{\rm poly}$, we do not adopt the RFFs and take the Tensor Sketch approximation [56] instead for simplicity.
Table 6 illustrates the comparisons on the detection performance among multiple alternative kernels. Actually, both the cosine-Laplacian kernel and the cosine-polynomial kernel cannot bring detection performance gains on top of the cosine kernel (CoP), which indicates that the $\ell_{1}$-distance relationship characterized by the Laplacian kernel and the inner-product information characterized by the polynomial kernel in the $\phi_{\rm cos}(\boldsymbol{z})$-space are less effective in promoting the separability between InD and OoD features. Thus, the cosine-Gaussian kernel is used and recommended in the proposed KPCA method for OoD detection.
Table 6: The detection results among a variety of kernels (ResNet50 trained on ImageNet-1K).
kernel OoD data sets AVERAGE
iNaturalist SUN Places Textures
[…]
PCA (no kernels) 95.46 52.01 97.98 44.86 97.99 45.19 46.22 87.77 84.41 57.46
Polynomial 96.03 53.07 98.26 42.84 97.85 45.02 95.50 47.96 96.91 47.22
Laplacian 94.65 50.25 94.68 50.29 95.28 49.80 94.66 50.34 94.82 50.17
Gaussian 94.46 50.83 95.17 50.33 94.80 50.46 95.09 50.80 94.88 50.60
Cosine (CoP) 67.25 83.41 75.53 79.93 82.48 73.83 8.33 98.29 58.40 83.86
Cosine-Polynomial 54.10 84.48 75.97 75.04 82.82 69.01 59.15 83.27 68.01 77.95
Cosine-Laplacian 76.18 77.95 77.54 76.70 84.47 70.16 11.97 97.57 62.54 80.60
Cosine-Gaussian (CoRP) 50.07 89.32 62.56 83.74 72.76 78.91 9.02 98.14 48.60 87.53
C.3 Sensitivity analysis
A comprehensive sensitivity analysis is executed to show the effects of hyper-parameters in CoP and CoRP. A common hyper-parameter in CoP and CoRP is the number of columns $q$ of the dimensionality-reduction matrix $\boldsymbol{U}^{\Phi}_{q}$. Additional hyper-parameters of CoRP include the bandwidth $\gamma$ of the Gaussian kernel and the number of RFFs $M$. In the following, we discuss the influence of each hyper-parameter, and report experiment results of the detection performance by varying one hyper-parameter with the others fixed.
@@ C.3 Sensitivity analysis
Figure 5: A sensitivity analysis on the explained variance ratio of CoP (top) and CoRP (bottom). The average FPR and AUROC values of OoD data sets in CIFAR10 and ImageNet-1K benchmarks are reported. The Gaussian kernel width $\gamma$ and the dimension $M$ of RFFs in CoRP are fixed.
Effect of $q$
$q$ indicates the number of preserved principal components and determines how much information captured by the subspace where the InD and OoD data is projected onto. $q$ is selected as the minimal number of principal components with the amount of information that exceeds the given explained variance ratio.
Figure 5 illustrates the detection performance of CoP and CoRP under varied explained variance ratios. On CIFAR10 and ImageNet-1K benchmarks, for CoP, a mild value of the explained variance ratio is suggested with around 90% for keeping the components. Regarding CoRP, a sufficiently large value of the explained variance ratio is no longer essential for CoRP on the ImageNet-1K benchmark, which might be due to that the 2 concatenated kernels make the useful information for distinguishing OoD samples more concentrated in less principal components.
@@
The Gaussian kernel width $\gamma$ directly affects the mapped data distribution. For a large $\gamma$, $k_{\rm gau}(\boldsymbol{z}_{1},\boldsymbol{z}_{2})=e^{-\gamma\left\|\boldsymbol{z}_{1}-\boldsymbol{z}_{2}\right\|_{2}^{2}}\approx 0$ for $\boldsymbol{z}_{1}\neq\boldsymbol{z}_{2}$, which indicates that the mapped features of $\boldsymbol{z}_{1}$ and $\boldsymbol{z}_{2}$ are (nearly) mutually-orthogonal. In this case, a PCA would become meaningless. For a small $\gamma$, the KPCA-based reconstruction errors will approach the standard PCA-based ones, shown by [57]. Figure 6 illustrates the detection FPR95 and AUROC values of CoRP w.r.t varied Gaussian kernel width $\gamma$ on CIFAR10 and ImageNet-1K benchmarks. Clearly, neither a too large nor a too small kernel width benefits the detection performance, and a mild value of $\gamma$ should be carefully tuned for different in-distribution data.
@@
The number of RFFs $M$ determines the approximation ability of RFFs towards the Gaussian kernel. As proved in [13], the larger the $M$, the better the RFFs approximate $k_{\rm gau}$. Figure 7 indicates the detection FPR95 and AUROC values of CoRP w.r.t. multiple values of the RFFs dimension $M$ on CIFAR10 and ImageNet-1K benchmarks. As $M$ increases, the detection performance gets improved since the RFFs better converge to the Gaussian kernel. Considering the computation efficiency of eigendecomposition on the covariance matrix of $\mathbb{R}^{M\times M}$, in the comparison experiments, we adopt $M=4m$ on CIFAR10 with $m=512$ for ResNet18, and $M=2m$ on ImageNet-1K with $m=2048$ for ResNet50 and $m=1280$ for MobileNet.
Figure 7: A sensitivity analysis on the dimension $M$ of RFFs of CoRP. The average detection FPR and AUROC values of OoD data sets in CIFAR10 and ImageNet-1K benchmarks are reported. The explained variance ratio and the Gaussian kernel width $\gamma$ are fixed.
Appendix D Supplementary theoretical results
@@
Recall $\boldsymbol{z}\in\mathbb{R}^{m}$ and suppose $\Phi:\mathbb{R}^{m}\rightarrow\mathbb{R}^{M}$ and $\boldsymbol{U}^{\Phi}\in\mathbb{R}^{M\times M}$ is the eigenvector matrix of the covariance matrix of the training data with $\boldsymbol{U}^{\Phi}=\left[\boldsymbol{U}^{\Phi}_{q},\boldsymbol{U}^{\Phi}_{p}\right]$ and $q+p=M$. For the reconstruction error $e^{\Phi}(\boldsymbol{\hat{z}})$ of a new test sample $\boldsymbol{\hat{z}}\in\mathbb{R}^{m}$ in the mapped $\Phi(\boldsymbol{z})$-space, we have:
$\left(e^{\Phi}(\boldsymbol{\hat{z}})\right)^{2}$ $=\left\|(\Phi(\boldsymbol{\hat{z}})-\boldsymbol{\mu}^{\Phi})-\boldsymbol{U}^{\Phi}_{q}\boldsymbol{U}_{q}^{\Phi\top}(\Phi(\boldsymbol{\hat{z}})-\boldsymbol{\mu}^{\Phi})\right\|_{2}^{2}$ (18)